OpenStudy (anonymous):

which of the following points does not lie on the graph of the function g(x)=2^x? A.(0,1) B.(2,4) C.(1,0) D.(-1,1/2) E. all of the above points are on the graph. I got B.(2,4) Is that correct? Check my math please and Thank you.

OpenStudy (anonymous):

Here is a picture of your graph, another way to look at this question is when you plug in the x value does not equal the y value. So, for A if x=0 then 2^0=1 so that is on the line B 2^2=4 so that is on the line C 2^1=2, but they are saying it should be 0 so that is not on the line. D 2^-1=0.5, so that is on the line

OpenStudy (anonymous):

Thank you ^^